OpenStudy (anonymous):

simplify 2/x^2+8x+15 + 1/x^2+11x+30

OpenStudy (anonymous):

do you know how to factorise

OpenStudy (anonymous):

you can 1st factorise the denominators find your common denominator and then add

OpenStudy (dumbcow):

\[\frac{2}{x^{2}+8x+15} + \frac{1}{x^{2}+11x+30}\] \[=\frac{2}{(x+3)(x+5)} + \frac{1}{(x+6)(x+5)}\] \[=\frac{2(x+6)}{(x+6)(x+3)(x+5)} + \frac{(x+3)}{(x+6)(x+3)(x+5)}\] \[=\frac{2(x+6) +(x+3)}{(x+6)(x+3)(x+5)}\] \[=\frac{3x+15}{(x+6)(x+3)(x+5)}\] \[=\frac{3(x+5)}{(x+6)(x+3)(x+5)}\] \[=\frac{3}{(x+6)(x+3)}\]